# Active Diffusion Subsampling

### v1

@@ Abstract
Subsampling is commonly used to mitigate costs associated with data acquisition, such as time or energy requirements, motivating the development of algorithms for estimating the fully-sampled signal of interest ${\mathbf{x}}$ from partially observed measurements ${\mathbf{y}}$. In maximum-entropy sampling, one selects measurement locations that are expected to have the highest entropy, so as to minimize uncertainty about ${\mathbf{x}}$. This approach relies on an accurate model of the posterior distribution over future measurements, given the measurements observed so far. Recently, diffusion models have been shown to produce high-quality posterior samples of high-dimensional signals using guided diffusion. In this work, we propose Active Diffusion Subsampling (ADS), a method for performing active subsampling using guided diffusion in which the model tracks a distribution of beliefs over the true state of ${\mathbf{x}}$ throughout the reverse diffusion process, progressively decreasing its uncertainty by choosing to acquire measurements with maximum expected entropy, and ultimately generating the posterior distribution $p({\mathbf{x}}\mid{\mathbf{y}})$. ADS can be applied using pre-trained diffusion models for any subsampling rate, and does not require task-specific retraining – just the specification of a measurement model. Furthermore, the maximum entropy sampling policy employed by ADS is interpretable, enhancing transparency relative to existing methods using black-box policies. Experimentally, we show that ADS outperforms fixed sampling strategies, and study an application of ADS in Magnetic Resonance Imaging acceleration using the fastMRI dataset, finding that ADS performs competitively with supervised methods. Code available at https://active-diffusion-subsampling.github.io/.
@@ 1 Introduction
In recent years, diffusion models have defined the state of the art in inverse problem solving, particularly in the image domain, through novel posterior sampling methods such as Diffusion Posterior Sampling (DPS) (Chung et al., 2022) and Posterior Sampling with Latent Diffuion (PSLD) (Rout et al., 2024). These methods are often evaluated on inverse imaging problems, such as inpainting, which is akin to image subsampling. Typical benchmarks evaluate inpainting ability using naïve subsampling masks such randomly masked pixels, or a box mask in the center of the image (Rout et al., 2024). In real-world applications, however, more sophisticated subsampling strategies are typically employed, for example, in Magnetic Resonance Imaging (MRI) acceleration (Lustig and Pauly, 2010, Bridson, 2007). These subsampling strategies are usually designed by domain experts, and are therefore not generalizable across tasks. Some recent literature has explored learning subsampling masks for various tasks (Bahadir et al., 2020, Baldassarre et al., 2016, Huijben et al., 2020, Van Gorp et al., 2021), but these methods typically depend on black-box policy functions, and require task-specific training. In this work, we introduce Active Diffusion Subsampling (ADS), an algorithm for automatically designing task- and sample-adaptive subsampling masks using diffusion models, without the need for further training or fine-tuning (Figure 2). ADS uses a white-box policy function based on maximum entropy sampling (Caticha, 2021), in which the model chooses sampling locations that are expected to maximize the information gained about the reconstruction target. In order to implement this policy, ADS leverages quantities that are already computed during the reverse diffusion process, leading to minimal additional computation. Our main contributions are thus as follows:
@@
  • Experimental validation against baseline sampling strategies and via application to MRI acceleration.
@@ 2 Related Work
Methods aiming to select maximally informative measurements appear in many domains, spanning statistics, signal processing, and machine learning, but sharing foundations in information theory and Bayesian inference. Optimal Bayesian Experimental Design (Lindley, 1956) aims to determine which experiment will be most informative about some quantity of interest $\theta$ (Rainforth et al., 2024), typically parameters of a statistical model. Active learning (Houlsby et al., 2011) performs an analogous task in machine learning, aiming to identify which samples, if included in the training set, would lead to the greatest performance gain on the true data distribution. While our method focuses on subsampling high-dimensional signals, it could also be interpreted as a Bayesian regression solver for the forward problem ${\mathbf{y}}=f({\mathbf{x}})+{\mathbf{n}}$, in which ${\mathbf{x}}$ is now seen as parameters for a model $f$, and the active sampler seeks measurements ${\mathbf{y}}$ which minimize uncertainty about the parameters ${\mathbf{x}}$, as in Bayesian experimental design.
A popular approach to subsampling from signal processing is compressive sensing, in which sparse signals are reconstructed from measurements with sub-Nyquist sampling rates (Rani et al., 2018), and typically applied in imaging and communication. Measurement matrices relating observed measurements to the signal of interest ${\mathbf{x}}$ are then designed so as to minimize reconstruction error on ${\mathbf{x}}$. In Bayesian compressive sensing, measurement matrices are designed so as to minimize a measure of uncertainty about the value of ${\mathbf{x}}$. Adaptive approaches, such as that of Braun et al. (2015), aim to decrease this uncertainty iteratively, by greedily choosing measurements that maximise the mutual information between ${\mathbf{x}}$ and ${\mathbf{y}}={\mathbf{U}}{\mathbf{x}}+{\mathbf{n}}$. These methods typically assume a Gaussian prior on ${\mathbf{x}}$, however, limiting the degree to which more complex prior structure can be used.
@@
More recently, a number of methods using deep learning to design subsampling strategies have emerged. These approaches typically learn subsampling strategies from data that minimize reconstruction error between ${\mathbf{x}}$ and ${\mathbf{y}}$. Methods by Huijben et al. (2020) and Bahadir et al. (2020) learn fixed sampling strategies, in which a single mask is designed a priori for a given domain, and applied to all samples for inference. These methods can be effective, but suffer in cases where optimal masks differ across samples. Sample-adaptive methods (Van Gorp et al., 2021, Bakker et al., 2020, Yin et al., 2021, Stevens et al., 2022) move past this limitation by designing sampling strategies at inference time. A popular application of such methods is MRI acceleration, spurred by the fastMRI benchmark (Zbontar et al., 2018), in which a full MRI image must be reconstructed from sub-sampled $\kappa$-space measurements. In A-DPS (Van Gorp et al., 2021), for example, a neural network is trained to build an acquisition mask consisting of $M$ $\kappa$-space lines iteratively, adaptively adding new lines based on the current reconstruction and prior context. Bakker et al. (2020) implements the same procedure using a reinforcement learning agent. One drawback of these methods is their reliance on black-box policies, making it difficult to detect and interpret failure cases. Generative approaches with transparent sampling policies circumvent this issue. For example, CLUDAS (Sanchez et al., 2020) is a generative approach to adaptive MRI acquisition design, using a conditional Wasserstein GAN to generate posterior samples over MRI image, and maximum-variance sampling in the $\kappa$-space as the measurement selection policy. The performance of generative approaches such as CLUDAS depends on how well they can model the true posterior distribution over ${\mathbf{x}}$ given observations. For this reason, ADS is designed to leverage state-of-the-art posterior sampling methods enabled by diffusion models, such as Posterior Sampling with Latent Diffusion (Rout et al., 2024). Furthermore, Diffusion Models have shown excellent performance in diverse domain, such as computer vision Dhariwal and Nichol (2021), medical imaging (Chung and Ye, 2022, Stevens et al., 2024), and natural language processing (Yu et al., 2022), and more, enhancing the potential for generalizability.
@@ 3.1 Bayesian Optimal Experimental Design
In Bayesian Optimal Experimental Design (Rainforth et al., 2024) and Bayesian Active Learning (Houlsby et al., 2011), the objective is to choose the optimal design, or set of actions $A=A^{*}$, leading to new observations of a measurement variable ${\mathbf{y}}$ that will minimise uncertainty about a related quantity of interest ${\mathbf{x}}$, as measured by entropy $H$. It was shown by Lindley (1956) that this objective is equivalent to finding the actions $A$ that maximise the mutual information $I$ between ${\mathbf{x}}$ and ${\mathbf{y}}$, i.e. selecting actions leading to observations of ${\mathbf{y}}$ that will be most informative about ${\mathbf{x}}$:
@@
This objective is commonly optimized actively, wherein the design is created iteratively by choosing actions that maximise mutual information, considering past observations when selecting new actions (Rainforth et al., 2024). This active paradigm invites an agent-based framing, in which the agent’s goal is to minimize its own uncertainty, and beliefs over ${\mathbf{x}}$ are updated as new measurements of ${\mathbf{y}}$ are taken. The active design objective can be formulated as follows, where $a_{t}$ is possible action at time $t$, $A_{t}=A_{t-1}\cup a_{t}$ is the set of actions $\{a_{0},...,a_{t}\}$ taken so far at time $t$, and ${\mathbf{y}}_{t-1}$ is the set of partial observations of the measurement variable ${\mathbf{y}}$ until time $t-1$:
@@
We can interpret the agent’s behaviour from Equation 2 as trying to maximise marginal uncertainty about ${\mathbf{y}}_{t}$, while minimising model uncertainty about what value ${\mathbf{y}}_{t}$ should take on given a particular ${\mathbf{x}}$ (Houlsby et al., 2011). Active designs are typically preferred over fixed designs, in which a set of actions is chosen up-front, as opposed to being chosen progressively as measurements are acquired (Rainforth et al., 2024). While fixed designs may be more computationally efficient, they are less sample-specific, which can lead to lower information gain about ${\mathbf{x}}$. Finally, it is worth noting that this active optimization scheme, while greedy, has been shown to be near-optimal due to the submodularity of conditional entropy (Golovin and Krause, 2011).
@@ 3.2 Subsampling
Since we are interested in the adaptive design of these masks, we formulate their generation by ${\mathbf{m}}={\mathbf{U}}(A_{t})^{\top}{\mathbf{U}}(A_{t})$, where the measurement matrix is now a function of the actions $A_{t}=\{a_{0},...,a_{t}\}$ taken by the agent up to time $t$. The $i^{\text{th}}$ element of that mask is defined as follows:
@@
The measurement model in (3) can now be extended to an active setting via ${\mathbf{y}}_{t}={\mathbf{U}}(A_{t}){\mathbf{x}}+{\mathbf{n}}_{t}$. Note too that in some applications we have an additional forward model $f$, mapping from the data domain to the measurement domain, yielding ${\mathbf{y}}_{t}={\mathbf{U}}(A_{t})f({\mathbf{x}})+{\mathbf{n}}_{t}$.
@@ 3.3 Posterior Sampling with Diffusion Models
Denoising diffusion models learn to reverse a stochastic differential equation (SDE) that progressively noises samples ${\mathbf{x}}$ towards a standard Normal distribution Song et al. (2020b). The SDE defining the noising process is as follows:
@@
where ${\mathbf{x}}(0)\in\mathbb{R}^{d}$ is an initial clean sample, $\tau\in[0,T]$, $\beta(\tau)$ is the noise schedule, and $w$ is a standard Wiener process, and ${\mathbf{x}}(T)\sim\mathcal{N}(0,I)$. According the following equation from Anderson (1982), this SDE can be reversed once the score function $\nabla_{{\mathbf{x}}}\log p_{\tau}({\mathbf{x}})$ is known, where $\bar{w}$ is a standard Wiener process running backwards:
@@
Following the notation by Ho et al. (2020) and Chung et al. (2022), the discrete setting of the SDE is represented using $x_{\tau}={\mathbf{x}}(\tau T/N),\beta_{\tau}=\beta(\tau T/N),\alpha_{\tau}=1-\beta_{\tau},\bar{\alpha}_{\tau}=\prod_{s=1}^{\tau}\alpha_{s}$, where $N$ is the numbers of discretized segments. The diffusion model achieves the SDE reversal by learning the score function using a neural network parameterised by $\theta$, $s_{\theta}({\mathbf{x}}_{\tau},\tau)\simeq\nabla_{{\mathbf{x}}_{\tau}}\log p_{\tau}({\mathbf{x}}_{\tau})$.
The reverse diffusion process can be conditioned on a measurement ${\mathbf{y}}$ to produce samples from the posterior $p({\mathbf{x}}|{\mathbf{y}})$. This can be done with substitution of the conditional score function $\nabla_{{\mathbf{x}}_{\tau}}\log p_{\tau}({\mathbf{x}}_{\tau})$ in (6). The intractability of the noise-perturbed likelihood $\nabla_{{\mathbf{x}}_{\tau}}\log p_{\tau}({\mathbf{y}}|{\mathbf{x}}_{\tau})$ which follows from refactoring the posterior using Bayes’ rule has led to various approximate guidance schemes to compute these gradients with respect to a partially-noised sample ${\mathbf{x}}_{\tau}$ (Chung et al., 2022, Song et al., 2023, Rout et al., 2024). Most of these rely on Tweedie’s formula, which can be thought of as a one-step denoising process from $\tau\rightarrow 0$, denoted $\mathcal{D}_{\tau}(.)$, using our trained diffusion model to estimate the true fully-denoised sample ${\mathbf{x}}_{0}$ as follows:
$\hat{{\mathbf{x}}}_{0}=\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{x}}_{\tau}]=\mathcal{D}_{\tau}({\mathbf{x}}_{\tau})=\frac{1}{\sqrt{\bar{\alpha}_{\tau}}}({\mathbf{x}}_{\tau}+(1-\bar{\alpha}_{\tau})s_{\theta}({\mathbf{x}}_{\tau},\tau).$ (7)
Diffusion Posterior Sampling (DPS) uses (7) to approximate $\nabla_{{\mathbf{x}}_{\tau}}\log p({\mathbf{y}}|{\mathbf{x}}_{\tau})\approx\nabla_{{\mathbf{x}}_{\tau}}\log p({\mathbf{y}}|\hat{{\mathbf{x}}}_{0})$. In the case of active subsampling, this leads to guidance term $\nabla_{{\mathbf{x}}_{\tau}}||{\mathbf{y}}_{t}-U(A_{t})f(\hat{{\mathbf{x}}}_{0})||^{2}_{2}$ indicating the direction in which $x_{\tau}$ should step in order to be more consistent with ${\mathbf{y}}_{t}=U(A_{t})f({\mathbf{x}})+{\mathbf{n}}$. The conditional reverse diffusion process then alternates between standard reverse diffusion steps and guidance steps in order to generate samples from the posterior $p({\mathbf{x}}\mid{\mathbf{y}}_{t})$.

### v2

@@ Abstract
Subsampling is commonly used to mitigate costs associated with data acquisition, such as time or energy requirements, motivating the development of algorithms for estimating the fully-sampled signal of interest ${\mathbf{x}}$ from partially observed measurements ${\mathbf{y}}$. In maximum-entropy sampling, one selects measurement locations that are expected to have the highest entropy, so as to minimize uncertainty about ${\mathbf{x}}$. This approach relies on an accurate model of the posterior distribution over future measurements, given the measurements observed so far. Recently, diffusion models have been shown to produce high-quality posterior samples of high-dimensional signals using guided diffusion. In this work, we propose Active Diffusion Subsampling (ADS), a method for designing intelligent subsampling masks using guided diffusion in which the model tracks a distribution of beliefs over the true state of ${\mathbf{x}}$ throughout the reverse diffusion process, progressively decreasing its uncertainty by actively choosing to acquire measurements with maximum expected entropy, ultimately producing the posterior distribution $p({\mathbf{x}}\mid{\mathbf{y}})$. ADS can be applied using pre-trained diffusion models for any subsampling rate, and does not require task-specific retraining – just the specification of a measurement model. Furthermore, the maximum entropy sampling policy employed by ADS is interpretable, enhancing transparency relative to existing methods using black-box policies. Experimentally, we show that through designing informative subsampling masks, ADS significantly improves reconstruction quality compared to fixed sampling strategies on the MNIST and CelebA datasets, as measured by standard image quality metrics, including PSNR, SSIM, and LPIPS. Furthermore, on the task of Magnetic Resonance Imaging acceleration, we find that ADS performs competitively with existing supervised methods in reconstruction quality while using a more interpretable acquisition scheme design procedure. Code is available at https://active-diffusion-subsampling.github.io/.
@@ 1 Introduction
In recent years, diffusion models have defined the state of the art in inverse problem solving, particularly in the image domain, through novel posterior sampling methods such as Diffusion Posterior Sampling (DPS) (Chung et al., 2022) and Posterior Sampling with Latent Diffusion (PSLD) (Rout et al., 2024). These methods are often evaluated on inverse imaging problems, such as inpainting, which is akin to image subsampling. Typical benchmarks evaluate inpainting ability using naïve subsampling masks such as randomly masked pixels, or a box mask in the center of the image (Rout et al., 2024). In real-world applications, however, more sophisticated subsampling strategies are typically employed, for example, in Magnetic Resonance Imaging (MRI) acceleration (Lustig & Pauly, 2010; Bridson, 2007). These subsampling strategies are usually designed by domain experts, and are therefore not generalizable across tasks. Some recent literature has explored learning subsampling masks for various tasks (Bahadir et al., 2020; Baldassarre et al., 2016; Huijben et al., 2020; Van Gorp et al., 2021), but these methods typically depend on black-box policy functions, and require task-specific training. In this work, we introduce Active Diffusion Subsampling (ADS), an algorithm for automatically designing task- and sample-adaptive subsampling masks using diffusion models, without the need for further training or fine-tuning (Figure 2). ADS uses a white-box policy function based on maximum entropy sampling (Caticha, 2021), in which the model chooses sampling locations that are expected to maximize the information gained about the reconstruction target. In order to implement this policy, ADS leverages quantities that are already computed during the reverse diffusion process, leading to minimal additional inference time. We anticipate that ADS can be employed by practitioners in various domains that use diffusion models for subsampling tasks but currently rely on non-adaptive subsampling strategies. Such domains include medical imaging, such as MRI, CT, and ultrasound (Shaul et al., 2020; Pérez et al., 2020; Faltin et al., 2011), geophysics (Campman et al., 2017; Cao et al., 2011), and astronomy (Feng et al., 2023). Our main contributions are thus as follows:
@@
  • Significant improvement in reconstruction quality relative to baseline sampling strategies.
  • Application to MRI acceleration and ultrasound scan-line subsampling.
@@ 2 Related Work
Methods aiming to select maximally informative measurements appear in many domains, spanning statistics, signal processing, and machine learning, but sharing foundations in information theory and Bayesian inference. Optimal Bayesian Experimental Design (Lindley, 1956) aims to determine which experiment will be most informative about some quantity of interest $\theta$ (Rainforth et al., 2024), typically parameters of a statistical model. Active learning (Houlsby et al., 2011) performs an analogous task in machine learning, aiming to identify which samples, if included in the training set, would lead to the greatest performance gain on the true data distribution. While our method focuses on subsampling high-dimensional signals, it could also be interpreted as a Bayesian regression solver for the forward problem ${\mathbf{y}}=f({\mathbf{x}})+{\mathbf{n}}$, in which ${\mathbf{x}}$ is now seen as parameters for a model $f$, and the active sampler seeks measurements ${\mathbf{y}}$ which minimize uncertainty about the parameters ${\mathbf{x}}$, relating it to the task of Bayesian experimental design.
@@
From a signal processing perspective, ADS can be characterized as a novel approach to adaptive compressive sensing, in which sparse signals are reconstructed from measurements with sub-Nyquist sampling rates (Rani et al., 2018), and typically applied in imaging and communication. Measurement matrices relating observed measurements to the signal of interest ${\mathbf{x}}$ are then designed so as to minimize reconstruction error on ${\mathbf{x}}$. In Bayesian compressive sensing, measurement matrices are designed so as to minimize a measure of uncertainty about the value of ${\mathbf{x}}$. Adaptive approaches, such as that of Braun et al. (2015), aim to decrease this uncertainty iteratively, by greedily choosing measurements that maximize the mutual information between ${\mathbf{x}}$ and ${\mathbf{y}}={\mathbf{U}}{\mathbf{x}}+{\mathbf{n}}$. These methods typically assume a Gaussian prior on ${\mathbf{x}}$, however, limiting the degree to which more complex prior structure can be used. More recently, a number of methods using deep learning to design subsampling strategies have emerged. These approaches typically learn subsampling strategies from data that minimize reconstruction error between ${\mathbf{x}}$ and ${\mathbf{y}}$. Methods by Huijben et al. (2020) and Bahadir et al. (2020) learn fixed sampling strategies, in which a single mask is designed a priori for a given domain, and applied to all samples for inference. These methods can be effective, but suffer in cases where optimal masks differ across samples. Sample-adaptive methods (Van Gorp et al., 2021; Bakker et al., 2020; Yin et al., 2021; Stevens et al., 2022) move past this limitation by designing sampling strategies at inference time. A popular application of such methods is MRI acceleration, spurred by the fastMRI benchmark (Zbontar et al., 2018), in which a full MRI image must be reconstructed from sub-sampled $\kappa$-space measurements. In A-DPS (Van Gorp et al., 2021), for example, a neural network is trained to build an acquisition mask consisting of $M$ $\kappa$-space lines iteratively, adaptively adding new lines based on the current reconstruction and prior context. Bakker et al. (2020) implements the same procedure using a reinforcement learning agent. One drawback of these methods is their reliance on black-box policies, making it difficult to detect and interpret failure cases. Generative approaches with transparent sampling policies circumvent this issue. For example, methods by Sanchez et al. (2020) and van de Camp et al. (2023) take a generative approach to adaptive MRI acquisition design, using variants of generative adversarial networks (GANs) to generate posterior samples over MRI images, and maximum-variance sampling in the $\kappa$-space as the measurement selection policy. The performance of such generative approaches depends on how well they can model the true posterior distribution over ${\mathbf{x}}$ given observations. This motivates our choice of diffusion models for generative modeling, as they have shown excellent performance in diverse domains, such as computer vision (Dhariwal & Nichol, 2021), medical imaging (Chung & Ye, 2022; Stevens et al., 2024), and natural language processing (Yu et al., 2022). The existing works discussed above have in common that they run the reconstruction model $N$ times for $N$ measurements during inference. Due to the slower iterative inference procedure used by diffusion models, running inference $N$ times may not be feasible in real-world applications, which are typically time-bound by the acquisition process. ADS alleviates this issue by using the Tweedie estimate of the posterior distribution as an approximate posterior, allowing $N$ measurements to be acquired in a single reverse diffusion process. We also highlight a related concurrent work by Elata et al. (2025), which uses denoising diffusion restoration models (DDRMs) to generate sensing matrices for adaptive compressed sensing, which differs from ADS in that it requires full DDRM inference in order to acquire each sample.
@@ 3.1 Bayesian Optimal Experimental Design
In Bayesian Optimal Experimental Design (Rainforth et al., 2024) and Bayesian Active Learning (Houlsby et al., 2011), the objective is to choose the optimal design, or set of actions $A=A^{*}$, leading to new observations of a measurement variable ${\mathbf{y}}$ that will minimize uncertainty about a related quantity of interest ${\mathbf{x}}$, as measured by entropy $H$. It was shown by Lindley (1956) that this objective is equivalent to finding the actions $A$ that maximize the mutual information $I$ between ${\mathbf{x}}$ and ${\mathbf{y}}$, i.e. selecting actions leading to observations of ${\mathbf{y}}$ that will be most informative about ${\mathbf{x}}$:
@@
This objective is commonly optimized actively, wherein the design is created iteratively by choosing actions that maximize mutual information, considering past observations when selecting new actions (Rainforth et al., 2024). This active paradigm invites an agent-based framing, in which the agent’s goal is to minimize its own uncertainty, and beliefs over ${\mathbf{x}}$ are updated as new measurements of ${\mathbf{y}}$ are taken. The active design objective can be formulated as follows, where $a_{t}$ is possible action at time $t$, $A_{t}=A_{t-1}\cup a_{t}$ is the set of actions $\{a_{0},...,a_{t}\}$ taken so far at time $t$, and ${\mathbf{y}}_{t-1}$ is the set of partial observations of the measurement variable ${\mathbf{y}}$ until time $t-1$:
@@
We can interpret the agent’s behavior from Equation 2 as trying to maximize marginal uncertainty about ${\mathbf{y}}_{t}$, while minimizing model uncertainty about what value ${\mathbf{y}}_{t}$ should take on given a particular ${\mathbf{x}}$ (Houlsby et al., 2011). Active designs are typically preferred over fixed designs, in which a set of actions is chosen up-front, as opposed to being chosen progressively as measurements are acquired (Rainforth et al., 2024). While fixed designs may be more computationally efficient, they are less sample-specific, which can lead to lower information gain about ${\mathbf{x}}$. Finally, it is worth noting that this active optimization scheme, while greedy, has been shown to be near-optimal due to the submodularity of conditional entropy (Golovin & Krause, 2011).
@@ 3.2 Subsampling
Since we are interested in the adaptive design of these masks, we express their generation as ${\mathbf{m}}={\mathbf{U}}(A_{t})^{\top}{\mathbf{U}}(A_{t})$, where the measurement matrix is now a function of the actions $A_{t}=\{a_{0},...,a_{t}\}$ taken by the agent up to time $t$. The $i^{\text{th}}$ element of that mask is defined as follows:
@@
The measurement model in equation 3 can now be extended to an active setting via ${\mathbf{y}}_{t}={\mathbf{U}}(A_{t}){\mathbf{x}}+{\mathbf{n}}_{t}$. Note too that in some applications we have an additional forward model $f$, mapping from the data domain to the measurement domain, yielding ${\mathbf{y}}_{t}={\mathbf{U}}(A_{t})f({\mathbf{x}})+{\mathbf{n}}_{t}$.
@@ 3.3 Posterior Sampling with Diffusion Models
Denoising diffusion models learn to reverse a stochastic differential equation (SDE) that progressively noises samples ${\mathbf{x}}$ towards a standard Normal distribution (Song et al., 2020). The (variance preserving) SDE defining the noising process is as follows:
@@
where ${\mathbf{x}}(0)\in\mathbb{R}^{d}$ is an initial clean sample, $\tau\in[0,T]$, $\beta(\tau)$ is the noise schedule, and $w$ is a standard Wiener process, and ${\mathbf{x}}(T)\sim\mathcal{N}(0,I)$. According to the following equation from Anderson (1982), this SDE can be reversed once the score function $\nabla_{{\mathbf{x}}}\log p_{\tau}({\mathbf{x}})$ is known, where $\bar{w}$ is a standard Wiener process running backwards:
@@
Following the notation by Ho et al. (2020) and Chung et al. (2022), the discrete setting of the SDE is represented using $x_{\tau}={\mathbf{x}}(\tau T/N),\beta_{\tau}=\beta(\tau T/N),\alpha_{\tau}=1-\beta_{\tau},\bar{\alpha}_{\tau}=\prod_{s=1}^{\tau}\alpha_{s}$, where $N$ is the numbers of discretized segments. The diffusion model achieves the SDE reversal by learning the score function using a neural network parameterized by $\theta$, $s_{\theta}({\mathbf{x}}_{\tau},\tau)\simeq\nabla_{{\mathbf{x}}_{\tau}}\log p_{\tau}({\mathbf{x}}_{\tau})$.
The reverse diffusion process can be conditioned on a measurement ${\mathbf{y}}$ to produce samples from the posterior $p({\mathbf{x}}|{\mathbf{y}})$. This can be done with substitution of the conditional score function $\nabla_{{\mathbf{x}}_{\tau}}\log p_{\tau}({\mathbf{x}}_{\tau})$ in equation 6. The intractability of the noise-perturbed likelihood $\nabla_{{\mathbf{x}}_{\tau}}\log p_{\tau}({\mathbf{y}}|{\mathbf{x}}_{\tau})$ which follows from refactoring the posterior using Bayes’ rule has led to various approximate guidance schemes to compute these gradients with respect to a partially-noised sample ${\mathbf{x}}_{\tau}$ (Chung et al., 2022; Song et al., 2023; Rout et al., 2024). Most of these rely on Tweedie’s formula (Efron, 2011; Robbins, 1992), which serves as a one-step denoising process from $\tau\rightarrow 0$, denoted $\mathcal{D}_{\tau}(.)$, yielding the Minimum Mean-Squared Error denoising of ${\mathbf{x}}_{\tau}$ (Milanfar & Delbracio, 2024). Using our trained score function $s_{\theta}({\mathbf{x}}_{\tau},\tau)$ we can approximate ${\mathbf{x}}_{0}$ as follows:
$\hat{{\mathbf{x}}}_{0}=\mathbb{E}[{\mathbf{x}}_{0}|{\mathbf{x}}_{\tau}]=\mathcal{D}_{\tau}({\mathbf{x}}_{\tau})=\frac{1}{\sqrt{\bar{\alpha}_{\tau}}}({\mathbf{x}}_{\tau}+(1-\bar{\alpha}_{\tau})s_{\theta}({\mathbf{x}}_{\tau},\tau))$ (7)
Diffusion Posterior Sampling (DPS) uses equation 7 to approximate $\nabla_{{\mathbf{x}}_{\tau}}\log p({\mathbf{y}}|{\mathbf{x}}_{\tau})\approx\nabla_{{\mathbf{x}}_{\tau}}\log p({\mathbf{y}}|\hat{{\mathbf{x}}}_{0})$. In the case of active subsampling, this leads to guidance term $\nabla_{{\mathbf{x}}_{\tau}}||{\mathbf{y}}_{t}-{\mathbf{U}}(A_{t})f(\hat{{\mathbf{x}}}_{0})||^{2}_{2}$ indicating the direction in which ${\mathbf{x}}_{\tau}$ should step in order to be more consistent with ${\mathbf{y}}_{t}={\mathbf{U}}(A_{t})f({\mathbf{x}})+{\mathbf{n}}$. The conditional reverse diffusion process then alternates between standard reverse diffusion steps and guidance steps in order to generate samples from the posterior $p({\mathbf{x}}\mid{\mathbf{y}}_{t})$. Finally, we note that a number of posterior sampling methods for diffusion models have recently emerged, and refer the interested reader to the survey by Daras et al. (2024).